The function f(x)= [x]- x is not integrable in [ 0,3] , where [ x] denote greatest integer function.


True or false with full explanation

1
Expert's answer
2022-02-24T04:07:01-0500

ANSWER. This is false.

EXPLANATION.

Since "[x]=\left\{\begin{matrix}
 0,&0\leq x <1\\ 
 1,&1\leq x <2\\ 
 2,&2\leq x <3\\ 
 3,& x=3
\end{matrix}\right." , then "f(x) =[x]-x=\left\{\begin{matrix}
 -x,&0\leq x <1\\ 
 1-x,&1\leq x <2\\ 
 2-x,&2\leq x <3\\ 
 0,& x=3
\end{matrix}\right." .



For all "x\in[0,3]" : "|f(x)|\leq 1" . Thus "f" is a bounded piecewise continuous function (continuous and monotonic in each interval "(k-1,k) k=1,2,3" ). Therefore "f" is integrable on "[0,3]" .